A waiter poured 2 gallons of coffee in 2 hours and
6 gallons of coffee in 6 hours. He also poured
1.5 gallons of tea in 3 hours and 2.5 gallons of tea
in 5 hours. Which drink did the waiter pour at a
greater rate?
1
Expert's answer
2020-10-12T13:18:06-0400

Rate of pouring coffee "= \frac{2}{2} = 1" gallon/hour

OR

Rate of pouring coffee "= \frac{6}{6} = 1" gallon / hour

And

Rate of pouring tea "= \frac{1.5}{3} = 0.5" gallon / hour

OR

Rate of pouring tea "= \frac{2.5}{5} = 0.5" gallon/ hour

As we know ,

"1 > 0.5"

So rate of pouring coffee is greater than rate of pouring tea.